Write the equation of the conic section x² + y² = 4 after a rotation of 45° about the origin.
(x')² - 2x'y' + (y')² = 4
(x')² + (y')² = 16
(x')² + 2x'y' + (y')² = 16
(x')² + (y')² = 4

Answer by ikleyn(52781)

.
Write the equation of the conic section x² + y² = 4 after a rotation of 45° about the origin.

(x')² - 2x'y' + (y')² = 4
(x')² + (y')² = 16
(x')² + 2x'y' + (y')² = 16
(x')² + (y')² = 4                <--- This and only this